OpenStudy (anonymous):

When answering questions 1–6, please separate your answers with a semicolon (;). For questions 1–6 determine whether each series is arithmetic or geometric. Then evaluate the series to the given term. 1. 1 + 3 + 9 + ...; S8 2. 5 + 7 + 9 + ...; S8 3. 25 + 32 + 39 + ...; S12 4. 256 + 192 + 144 + ...; S7 5. 1 – 3 + 9 – 27 + ...; S10 6. –14 + (–7) + 0 + 7 + ...; S10

OpenStudy (anonymous):

@Ket-kat this is the one

OpenStudy (anonymous):

1. + 3 + 9 terms .............upto 8 The given series is geometric series like a, ar, ar^2, ---------- where a is first term and r is common ratio. The sum of n terms in geometric series is given by, Sn = a(r^n - 1)/(r - 1) Here a = 1, r = 3 and n = 8 S8 = 1(3^8 - 1)/(3-1) = (6561 - 1)/2 = 3280

OpenStudy (anonymous):

2. arithmetic 96

OpenStudy (anonymous):

3.arithmetic add by 7

OpenStudy (anonymous):

4.geometric 5.geometric 6.arithmetic

OpenStudy (anonymous):

@Ket-kat keep going and i'll give you a medal(: thanks this really helps!

OpenStudy (anonymous):

@Ket-kat Thanks (: but u forgot to evaluate the series to the given term for some of them